OpenStudy (anonymous):

Suppose an ISP has the addres like 101.101.128/17.He want to create 4 subnets.What are the prefix (a.b.c.d/x) for this four subnets?

OpenStudy (anonymous):

@izanagi.wielder

OpenStudy (anonymous):

wait i will have think

OpenStudy (anonymous):

i forgot to mention every block mus be the same number of IP address

OpenStudy (anonymous):

101.101.101.64/28 101.101.101.80/28 101.101.101.96/28 101.101.101.112/28

OpenStudy (anonymous):

how u get it? izanagi

OpenStudy (anonymous):

well my bro is a cs learner i asked him

OpenStudy (anonymous):

lol,ask for him how to get the result please

OpenStudy (anonymous):

you can contact him later @Ishaan94

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

101.101.128.0/19 101.101.160.0/19 101.101.192.0/19 101.101.224.0/19

OpenStudy (anonymous):

these are other answers my bro says break it into 4 equal pieces

OpenStudy (anonymous):

i do it like this too,but the ans is equal to first form,and i dont know why we use /28 instead /19

OpenStudy (anonymous):

to make four subnets i must borrow 2 bits from the host,i think

OpenStudy (anonymous):

i am not that deep in it you better contact him

OpenStudy (anonymous):

ok thanks izanagi

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

You need to convert to binary, each is 8 long in binary so 101.101.128.0/17 converts to: 01100101.01100101.10000000.00000000 The 17 means the 17 first bits will be equal to each adress, this equals to: 01100101.01100101.1 (8+8+1) Now you have 0000000.00000000 to make 4 equal subnets with, 15 bits To get 4 different we need to use 2 bits, 00 / 01 / 10 / 11 This will give us the 4 subnets, now with added 2 bits: 01100101.01100101.100 = 101.101.128.0/19 01100101.01100101.101 = 101.101.160.0/19 01100101.01100101.110 = 101.101.192.0/19 01100101.01100101.111 = 101.101.224.0/19 Hope this helps :)

OpenStudy (anonymous):

@SigurdS i thougth samething but the answ in book is: 101.101.101.64/28 101.101.101.80/28 101.101.101.96/28 101.101.101.112/28

OpenStudy (anonymous):

I think the book is wrong, ask your professor but those are not the same subnets as the one mentioned in the question, as the extra 101 block will not fall into the 128, (01100101 vs 10000000) Books often do silly mistakes like that! (I am not a 100% expert on subnetting so I might be missing something)

OpenStudy (anonymous):

thanks